OpenStudy (anonymous):

i need a quadratic equation to: narrower 4, up3, right 2..

OpenStudy (turingtest):

I don't understand... what math is this?

OpenStudy (anonymous):

algebra 2

OpenStudy (turingtest):

I am not familiar with the terms "narrower, up, and right" as common. Are they specific to some graph (ellipse, hyperbola, etc.). If not please explain, I doubt many here know them.

OpenStudy (anonymous):

narrow is what shape its going to be in the graph, up and right is where you have the intercepts

OpenStudy (saifoo.khan):

Can u draw it?

OpenStudy (anonymous):

for a project you have to make a graph, Narrower: 4 goes up 3 and 2 to the right, how do you write that in a absolute value equation? i dont know how to make that in a equation or how to graph it.

OpenStudy (turingtest):

These terms are unfamiliar to most of us here. I agree with saifoo that if you can draw the graph that would help most, and pointing out the coordintes (i.e. what you mean by these terms) would help greatly.

OpenStudy (anonymous):

(2,3) the shape is narrow , i need a absloute value equation

OpenStudy (anonymous):

Its an equation for f(x)

OpenStudy (turingtest):

3=2^2+n--->n=-1 y=x^2-1sounds like a possibility, but your question is too unclear to be sure this is what you mean. Since you said you need absolute value you could write\[y=\left| x^2 \right|-1\] which changes nothing.